# Do Comments and Expertise Still Matter? An Experiment on Programmers’ Adoption of AI-Generated JavaScript Code

### v1

@@ Abstract
This paper investigates the factors influencing programmers’ adoption of AI-generated JavaScript code recommendations. It extends prior research by (1) utilizing objective (as opposed to the typically self-reported) measurements for programmers’ adoption of AI-generated code and (2) examining whether AI-generated comments added to code recommendations and development expertise drive AI-generated code adoption. We tested these potential drivers in an online experiment with 173 programmers. Participants were asked to answer some questions to demonstrate their level of development expertise. Then, they were asked to solve a LeetCode problem without AI support. After attempting to solve the problem on their own, they received an AI-generated solution to assist them in refining their solutions. The solutions provided were manipulated to include or exclude AI-generated comments (a between-subjects factor). Programmers’ adoption of AI-generated code was gauged by code similarity between AI-generated solutions and participants’ submitted solutions, providing a more reliable and objective measurement of code adoption behaviors. Our findings revealed that the presence of comments significantly influences programmers’ adoption of AI-generated code regardless of the participants’ development expertise.
@@ 1 Introduction
Our methodology not only deepens our understanding of the determinants of AI code adoption but also ensures the reliability of our findings through more objective measures.
@@ 2.3 Development Expertise
While previous studies have not thoroughly evaluated how development expertise influences programmers’ adoption of AI-generated code, related research on its impact on code comprehension offers a valuable starting point. Development expertise comprises multiple aspects, including general programming experience, programming experience in particular languages, familiarity with task-specific knowledge, and personality traits [22]. Previous studies have explored the influence of development expertise on code comprehension [23, 24]. Lawrie et al. discovered that increased years of work experience and schooling resulted in a better capability of comprehending source code [23]. Wagner and Wyrich studied the influence of intelligence and personality traits on code comprehension [24]. The results of their study showed that there was a strong positive correlation between participants’ general intelligence and their performance on code comprehension and a weaker negative correlation between participants’ conscientiousness and code comprehension [24].
@@ 3.1 Research Questions
Previous research has revealed the pivotal role of comments in code comprehension [18]. We hypothesize that comments still matter in programmers’ adoption of AI-generated code. We assume that the presence of comments in AI-generated code would improve programmers’ comprehension of the code, thereby leading to higher rates of code adoption. Our formal hypothesis for RQ1 is:
@@
$\mathbf{RQ_{2}}$. How does ($B_{1}$) development expertise influence programmers’ adoption of AI-generated code?
Programmers with richer development expertise tend to be more experienced in programming and more familiar with task-specific knowledge [22], which could potentially enable them to have a better understanding of AI-generated code without comments. We hypothesize that the influence of ($A_{1}$) presence of comments on programmers’ adoption of AI-generated code would be moderated by ($B_{1}$) development expertise, such that the positive effect of comments on adoption decreases as development expertise increases. The formal hypothesis for RQ2 is:
@@ 3.2.1 Experiment Structure
To ensure the reliability and objectivity of our findings, we designed a survey-based experiment with a structured approach. By having participants first solve a programming problem on their own, we create a baseline for their understanding and approach. Then, by introducing an AI-generated solution with or without comments, we can objectively measure how these factors influence their adoption behaviors. This methodology allows us to isolate the effects of comments and development expertise, providing a clear understanding of how these specific factors impact programmers’ willingness to adopt AI-generated code. Additionally, this design discourages participants from passively accepting AI-generated solutions, prompting them to critically evaluate and potentially integrate them with their own coding efforts. Conducting the experiment online with a consistent environment helps maintain control over external variables, ensuring that our results are robust and generalizable.
@@
The second page collected participants’ demographics to obtain the raw data of the (B) programmer factors, including ($B_{11}$) programming experience, ($B_{12}$) programming language proficiency, and ($B_{4}$) attitudes toward AI code generators. One of the programmer factors, ($B_{13}$) domain knowledge familiarity, would be investigated on the fifth page rather than the second page. Domain knowledge is the knowledge utilized to solve the programming problem on the third and fourth pages. We did not present questions about domain knowledge familiarity on page two to avoid potential bias. Asking about domain knowledge before presenting the programming problem might give participants hints, which could affect their responses. For instance, if solving a programming problem required knowledge of the data structure tree, then the domain knowledge familiarity question would ask the participants how familiar they were with the data structure tree. Asking this question would give participants hints that solving the programming problem required using the data structure tree, which could unintentionally make it easier for the participants to understand the provided snippets.
@@
The screenshot of the third page is shown in Fig. 2. To provide participants with a coding experience similar to Microsoft Virtual Studio Code (VSCode) [27], we integrated Monaco Editor, a code editor that empowers VSCode [28], into the survey website. Our code editor facilitated syntax highlighting and auto-completion features akin to VSCode, enabling participants to quickly adapt to the coding environment in our survey. Participants could execute their code with test cases specified by themselves or with predefined test cases. There were ten different predefined test cases for each programming problem, and these predefined test cases were hidden from the participants. Participants could view comprehensive execution results, such as error types, error messages, and the line number where an error occurs. A time limit was imposed for solving the programming problem on this page, ranging from a minimum of one minute to a maximum of ten minutes. The lower time limit was set to encourage participants to exert effort, while the upper time limit was employed to simulate real-life coding conditions, where time constraints are typically present when completing coding tasks and to enable scheduling of sessions.
@@
On the fourth page, we provided a code snippet and notified the participants that it was generated by an AI code generator. They were also notified that they could modify the code written by themselves on page three with the assistance of the provided AI-generated code. They had the option to use all the provided code, integrate part of the provided code into their original code, or choose not to use the provided code at all. We employed the same timer, code editor, and code execution features on page three.
@@ 3.2.2 Motivation and Engagement Mechanisms
The second motivation mechanism was full-screen mode, which encouraged participants to stay in full-screen mode while doing the survey. We motivated participants to stay in full-screen mode for two reasons. Firstly, staying in full-screen mode helps participants focus on the survey. Secondly, maintaining the survey in full-screen mode helps to prevent participants from seeking external assistance, such as information from the Internet and solutions created by the participants themselves using AI tools. Our goal was for participants to interact solely with the AI-generated solutions provided in our survey to gauge their adoption accurately. External information could influence their adoption of the provided AI-generated code in the survey, which might introduce bias.
@@ 3.3.3 AI-generated Solutions
We have created two versions of AI-generated solutions for each programming problem: solutions with comments and solutions without comments. Every survey was assigned one of these two versions. On the fourth page of the survey, participants used the assigned AI-generated solutions to assist in modifying the code written by themselves on page three. The provided AI-generated code snippets were generated by ChatGPT 4. To generate the solutions, we copied the problem description, including examples, into ChatGPT 4. Then, we added the prompt ”Please solve this problem in JavaScript” and included the template code from LeetCode, which provides function signatures. We then added a final prompt based on the version: ”Please do not add any comment” for the no-comment version and ”Please add meaningful comments” for the commented version. We confirmed that all the code snippets generated by ChatGPT in this research could be executed correctly and pass all test cases on LeetCode.
@@ 3.4.3 Dependent Variable
To measure the adoption of AI-generated code by programmers, we devised a novel objective metric. Participants were tasked with writing a code snippet with the assistance of an AI-generated solution (page four of the survey). The degree of adoption was then quantified by comparing the similarity between the code submitted by the participants and the AI-generated code provided on page four. This method provides an objective assessment of the code adoption behavior.
@@
The dimension of code similarity level showed the level at which the comparison was conducted. The possible values for the dimension of code similarity level were line level and token sequence level. The line-level code similarity metrics calculated the ratio of identical lines to total lines. Only lines that were exactly the same in both the AI-generated code and the participants’ submitted code would be counted as a match (differences in whitespace and comments were ignored). Token-sequence-level code similarity metrics compute similarity by comparing the token sequences of two code snippets, a process carried out using Jplag [35]. The token is a basic element of the code abstraction [35, 36]. A token sequence represents some fractions of the code snippet without necessarily representing a line of code. Token sequence abstracts the code, providing a mapping to the code structure without being influenced by different identifier names and comments [36]. Some code similarity-checking tools were developed to detect plagiarism in students’ programming assignments at a token-sequence level [37], including MOSS [36] and JPlag [35]. We used JPlag [35] to calculate the code similarity at the token-sequence level  [^1] We utilized the JPlag Java package from https://github.com/jplag/JPlag to calculate the token-sequence-level similarity. Line-level code similarity investigated participants’ direct copy-and-paste behavior. In contrast, the token-sequence-level code similarity focused on the abstract structure of the code.
@@ 3.5.2 Post-Survey Screening Mechanism
Age Minimum: 18, Maximum: 68, Mean: 27, SD: 7.748; Age distribution: $<$23 (25%), 23–32 (50%), $>$32 (25%)
@@ 4.1.1 Development Expertise
We then performed PCA on $B_{11a}$, $B_{11b}$, $B_{12a}$, $B_{12b}$, $B_{13a}$, and $B_{13b}$ to compute ($A_{3}$) development expertise. We tested the linearity assumption [45] and the sampling adequacy assumption [46] to ensure that the collected data were suitable for PCA. We then used the scree plot criterion [45] to choose latent factors that represent the observed variables. The scree plot criterion [45] suggested that only the first factor should be chosen. Keeping the first component explained 54.729% of the variance, which was acceptable [45]. The component matrix is shown in Table 5, indicating that the chosen factor could cover all observed variables. The interpretation of the data was consistent with the concepts of development expertise, since it involves positive coefficients for overall programming experience, programming experience in particular programming languages, and task-specified knowledge familiarity.
Table 5: Component matrix for development expertise
Observed Variable $B_{11\text{a}}$ $B_{11\text{b}}$ $B_{12\text{a}}$ $B_{12\text{b}}$ $B_{13\text{a}}$ $B_{13\text{b}}$
Coefficient 0.751 0.813 0.748 0.758 0.722 0.635
@@ 4.1.2 Latent Code Similarity
To simplify the data analysis and see the big picture of code similarity, we used PCA to generate the independent variable latent code similarity inferred from the four code similarity metrics listed in Table LABEL:tab:similarity-metrics. The procedure of computing latent code similarity was the same as computing development expertise. One latent factor was selected to represent the TP, TA, LP, and LA. The retained factor explained 82.550% of the total variances. As shown in Table 6, the retained latent variable could capture all the features of the four different code similarity metrics.
Table 6: Component matrix for latent code similarity
Code Similarity Metric TP TA LP LA
Coefficient 0.915 0.905 0.896 0.918
We also observed the difference between participants’ code written on page three and page four. There were seven participants who did not modify their code after being provided with the AI-generated code. This result indicated that the majority of participants (95.95%) were influenced by the AI-generated solution.

### v2

@@ Abstract
This paper investigates the factors influencing programmers’ adoption of AI-generated JavaScript code recommendations within the context of lightweight, function-level programming tasks. It extends prior research by (1) utilizing objective (as opposed to the typically self-reported) measurements for programmers’ adoption of AI-generated code and (2) examining whether AI-generated comments added to code recommendations and development expertise drive AI-generated code adoption. We tested these potential drivers in an online experiment with 173 programmers. Participants were asked to answer some questions to demonstrate their level of development expertise. Then, they were asked to solve a LeetCode problem without AI support. After attempting to solve the problem on their own, they received an AI-generated solution to assist them in refining their solutions. The solutions provided were manipulated to include or exclude AI-generated comments (a between-subjects factor). Programmers’ adoption of AI-generated code was gauged by code similarity between AI-generated solutions and participants’ submitted solutions, providing a behavioral measurement of code adoption behaviors. Our findings revealed that, within the context of function-level programming tasks, the presence of comments significantly influences programmers’ adoption of AI-generated code regardless of the participants’ development expertise.
@@ 1 Introduction
Our methodology not only deepens our understanding of the determinants of AI code adoption but also provides objective measurements of code adoption. Objective data is not necessarily superior to subjective self-reported data, but it could provide valuable complements to prior research.
@@ 2.3 Development Expertise
While previous studies have not thoroughly evaluated how development expertise influences programmers’ adoption of AI-generated code, related research on its impact on code comprehension offers a valuable starting point. Development expertise comprises multiple aspects, including general programming experience, programming experience in particular languages, familiarity with task-specific knowledge, and personality traits [22]. Previous studies have explored the influence of development expertise on code comprehension [23, 24]. Lawrie et al. discovered that increased years of work experience and schooling resulted in a better capability for comprehending source code [23]. Wagner and Wyrich studied the influence of intelligence and personality traits on code comprehension [24]. The results of their study showed that there was a strong positive correlation between participants’ general intelligence and their performance on code comprehension and a weaker negative correlation between participants’ conscientiousness and code comprehension [24].
@@ 3.1 Research Questions
Our hypothesis was developed based on the Unified Theory of Acceptance and Use of Technology (UTAUT) [14]. Fig. 1 illustrates how the presence of comments shapes programmers’ adoption intentions and how developers’ expertise moderates the effect of comments on adoption intentions, with all relationships grounded in UTAUT. According to UTAUT, lower effort expectancy is linked to higher adoption intention [14]. Moreover, existing research has underscored the critical role of comments in supporting code comprehension [18]. Building on these theoretical and empirical foundations, we propose that the inclusion of comments in AI-generated code will reduce the time cost and cognitive burdens programmers incur to comprehend the code. This reduction aligns with UTAUT’s emphasis on effort expectancy, as lowered comprehension costs directly translate to decreased effort expectancy, ultimately influencing adoption intentions. Our formal hypothesis for RQ1 is:
@@
Figure 1: Hypothesis model based on UTAUT
$\mathbf{RQ_{2}}$. How does ($B_{1}$) development expertise moderate the influence of ($A_{1}$) presence of comments on programmers’ adoption of AI-generated code?
Programmers with richer development expertise tend to be more experienced in programming and more familiar with task-specific knowledge [22], which could potentially enable them to have a better understanding of AI-generated code without comments. Aligning with UTAUT, experienced programmers—equipped with stronger code comprehension abilities—perceive a smaller reduction in effort expectancy when comments are present. This is because the reductions in time costs and cognitive burdens facilitated by comments hold less significance for them, given their capacity to effectively comprehend code regardless of whether comments are included. We hypothesize that the influence of ($A_{1}$) presence of comments on programmers’ adoption of AI-generated code would be moderated by ($B_{1}$) development expertise, such that the positive effect of comments on adoption decreases as development expertise increases. The formal hypothesis for RQ2 is:
@@ 3.2.1 Experiment Structure
To ensure the reliability of our findings, we designed a survey-based experiment with a structured approach. By having participants first solve a programming problem on their own, we create a baseline for their understanding and approach. Then, by introducing an AI-generated solution with or without comments, we can objectively measure how these factors influence their adoption behaviors.
This methodology allows us to isolate the effects of comments and development expertise, providing a clear understanding of how these specific factors impact programmers’ willingness to adopt AI-generated code. Additionally, this design discourages participants from passively accepting AI-generated solutions, prompting them to critically evaluate and potentially integrate them with their own coding efforts. Conducting the experiment online with a consistent environment helps maintain control over external variables, ensuring that our results are robust and generalizable.
@@
The second page collected participants’ demographics to capture the raw data for programmer factors, including ($B_{11}$) programming experience, ($B_{12}$) programming language proficiency, and ($B_{4}$) attitudes toward AI code generators. We postponed measuring ($B_{13}$) domain knowledge familiarity until page five to avoid biasing participants. Domain knowledge is the knowledge utilized to solve the programming problem on the third and fourth pages. Asking about domain knowledge upfront could unintentionally reveal what concepts were relevant to the upcoming programming task, influencing participants’ responses. For example, if the task involved tree structures, asking about familiarity with trees in advance could hint at the solution and influence participants’ responses.
@@
The screenshot of the third page is shown in Fig. 3. As illustrated, the coding problem interface consists of several distinct regions, including the code editor region, test case region, timer region, and problem description region.
@@
To provide participants with a coding experience similar to Microsoft Visual Studio Code (VSCode) [27], we integrated Monaco Editor, the code editor behind VSCode [28], into the code editor region. This integration enabled features such as syntax highlighting and auto-completion, allowing participants to quickly adapt to the environment.
In the test case region, participants could execute their code using either custom test cases or predefined ones. Each programming problem included ten predefined test cases, which were hidden from participants. As shown in Fig. 4, the test case region also supported switching to an output console, where participants could view comprehensive execution results, including error types, error messages, and the line number where an error occurred.
Figure 4: Overview of the test case region
A time limit was imposed on this page, ranging from a minimum of one minute to a maximum of ten minutes. The lower bound encouraged participants to make a genuine effort, while the upper bound simulated real-life time constraints and helped maintain a manageable study schedule.
On the fourth page, we provided a code snippet and notified the participants that it was generated by an AI code generator. They were also notified that they could modify the code written by themselves on page three with the assistance of the provided AI-generated code. Participants were explicitly informed that they had the option to use all the provided code, integrate part of the provided code into their original code, or choose not to use the provided code at all. We employed the same timer, code editor, and code execution features on page three.
@@ 3.2.2 Motivation and Engagement Mechanisms
The second motivation mechanism was full-screen mode, which encouraged participants to stay in full-screen mode while doing the survey. We motivated participants to stay in full-screen mode for two reasons. Firstly, staying in full-screen mode helps participants focus on the survey. Secondly, maintaining the survey in full-screen mode helps to prevent participants from seeking external assistance, such as information from the Internet and solutions created by the participants themselves using AI tools. Our goal was that participants only interact with the AI-generated solutions provided in our survey to gauge their adoption accurately. External information could influence their adoption of the provided AI-generated code in the survey, which might introduce bias.
@@ 3.3.3 AI-generated Solutions
We have created two versions of AI-generated solutions for each programming problem: solutions with comments and solutions without comments. Every survey was assigned one of these two versions. On the fourth page of the survey, participants used the assigned AI-generated solutions to assist in modifying the code written by themselves on page three.
The provided AI-generated code snippets were generated by ChatGPT 4 Turbo. To generate the solutions, we copied the problem description, including examples, into ChatGPT 4 Turbo. Then, we added the prompt ”Please solve this problem in JavaScript” and included the template code from LeetCode, which provides function signatures. We then added a final prompt based on the version: ”Please do not add any comment” for the no-comment version and ”Please add meaningful comments” for the commented version. We confirmed that all the code snippets generated by ChatGPT in this research could be executed correctly and passed all test cases on LeetCode. We also verified that all comments in the AI-generated code were meaningful and contributed to understanding the solution, rather than being irrelevant or nonsensical. An example of the AI-generated solution with and without comments is shown in Fig. 6.
* @param {string} s
* @return {number}
var longestValidParentheses = function (s) {
let maxLen = 0;
let stack = [-1];
for (let i = 0; i < s.length; i++) {
if (s[i] === "(") {
stack.push(i);
} else {
stack.pop();
if (stack.length === 0) {
stack.push(i);
} else {
const currentLen = i - stack[stack.length - 1];
maxLen = Math.max(maxLen, currentLen);
return maxLen;
Listing 1: An example of the AI-generated code without comments
* @param {string} s
* @return {number}
var longestValidParentheses = function (s) {
let maxLen = 0;
// Use a stack to track the indices of unmatched parentheses.
// -1 handles edge cases where the substring starts at index 0
let stack = [-1];
for (let i = 0; i < s.length; i++) {
if (s[i] === "(") {
stack.push(i);
} else {
// Pop for matching ’)’
stack.pop();
if (stack.length === 0) {
// If stack is empty, it means this ’)’ is unmatched.
// Push its index to reset the base for future valid substrings.
stack.push(i);
} else {
// Calculate current valid substring length
const currentLen = i - stack[stack.length - 1];
maxLen = Math.max(maxLen, currentLen);
return maxLen;
Listing 2: An example of the AI-generated code with comments
Figure 6: An example of the AI-generated code with and without comments
@@ 3.4.3 Dependent Variable
Building on research on algorithm appreciation [35], which suggests that individuals often rely on algorithmic outputs even without consciously acknowledging their influence, we define code adoption as behavioral reliance on AI-generated code rather than verbal declaration. Consistent with prior research that favors observable behavior [35], we measure this reliance using code similarity metrics—that is, the extent to which participants incorporated AI-generated code into their final submissions.
To operationalize this definition, we devised a novel objective metric. Participants were tasked with writing a code snippet with the assistance of an AI-generated solution (page four of the survey). The degree of adoption was then quantified by comparing the similarity between the code submitted by the participants and the AI-generated code provided on page four. This method provides an objective assessment of the code adoption behavior.
@@
The dimension of code similarity level showed the level at which the comparison was conducted. The possible values for the dimension of code similarity level were line level and token sequence level. The line-level code similarity metrics calculated the ratio of identical lines to total lines. Only lines that were exactly the same in both the AI-generated code and the participants’ submitted code would be counted as a match (differences in whitespace and comments were ignored). Token-sequence-level code similarity metrics compute similarity by comparing the token sequences of two code snippets, a process carried out using Jplag [36]. The token is a basic element of the code abstraction [36, 37]. A token sequence represents some fractions of the code snippet without necessarily representing a line of code. Token sequence abstracts the code, providing a mapping to the code structure without being influenced by different identifier names and comments [37].
Some code similarity-checking tools were developed to detect plagiarism in students’ programming assignments at a token-sequence level [38], including MOSS [37] and JPlag [36]. We used JPlag [36] to calculate the code similarity at the token-sequence level [^1] We utilized the JPlag Java package from https://github.com/jplag/JPlag to calculate the token-sequence-level similarity. Line-level code similarity investigated participants’ direct copy-and-paste behavior. In contrast, the token-sequence-level code similarity focused on the abstract structure of the code.
JPlag captures code similarity at an abstract, structural level by detecting patterns such as iteration loops, control flow, and the use of data structures in the AI-generated code [36]. This enables JPlag to identify participants’ adoption of AI-generated algorithms even when the code is not copied verbatim but reimplemented with similar underlying logic. For example, in the case shown in Fig. 8, both the AI-generated code and the participant’s submission used recursion to validate a binary tree. Despite this shared algorithmic structure, the two code snippets appeared quite different on the surface, with only two exact line matches. The line-level similarity from the AI perspective (LA) was just 0.13, whereas the token-sequence-level similarity from the AI perspective (TA), as computed by JPlag, was 0.75. This example illustrates JPlag’s ability to capture structural similarity beyond superficial textual resemblance.
var isValidBST = function (root) {
function validate(node, min, max) {
// An empty tree is a valid BST
if (!node) return true;
// The node’s value must be within the min and max bounds
if (
(min !== null && node.val <= min) $|$
(max !== null && node.val >= max)
return false;
// Recursively check the left subtree and right subtree
// For the left subtree, update the max bound
// For the right subtree, update the min bound
return (
validate(node.left, min, node.val) && validate(node.right, node.val, max)
// Initially, there are no bounds on the root’s value
return validate(root, null, null);
Listing 3: An example of the AI-generated code
var isValidBST = function (root) {
return isValidNode(root, -Infinity, Infinity);
function isValidNode(node, min, max) {
if (node === null) return true;
const curValidate = node.val > min && node.val < max;
// recursively check left subtree
const leftValidate = isValidNode(node.left, min, node.val);
// recursively check right subtree
const rightValidate = isValidNode(node.right, node.val, max);
return curValidate && leftValidate && rightValidate;
Listing 4: An example of the user-submitted code
Figure 8: A comparison between AI-generated code and user-submitted code
@@ 3.5.2 Post-Survey Screening Mechanism
Age Minimum: 18, Maximum: 68, Mean: 27, SD: 7.748; Age distribution: <<23 (25%), 23–32 (50%), >>32 (25%)
@@ 4.1.1 Development Expertise
We then performed PCA on $B_{11a}$, $B_{11b}$, $B_{12a}$, $B_{12b}$, $B_{13a}$, and $B_{13b}$ to compute ($A_{3}$) development expertise. We tested the linearity asssumption and sampling adequacy assumption [46] to ensure that the collected data were suitable for PCA. We have tested the Kaiser-Meyer-Olkin (KMO) measure [46] and Bartlett’s test of sphericity [47] to ensure the sampling adequacy assumption is not violated. The detailed information of KMO measure, Bartlett’s test of sphericity, and other detail of PCA analysis is shown in A. We then used the scree plot criterion [48] to choose latent factors that represent the observed variables. The scree plot criterion [48] suggested that only the first factor should be chosen. Keeping the first component explained 54.729% of the variance, which was acceptable [48]. The component matrix is shown in Table 11, indicating that the chosen factor could cover all observed variables. The interpretation of the data was consistent with the concepts of development expertise, since it involves positive coefficients for overall programming experience, programming experience in particular programming languages, and task-specified knowledge familiarity.
@@ 4.1.2 Latent Code Similarity
To simplify the data analysis and see the big picture of code similarity, we used PCA to generate the independent variable latent code similarity inferred from the four code similarity metrics listed in Table LABEL:tab:similarity-metrics. The procedure of computing latent code similarity was the same as computing development expertise. One latent factor was selected to represent the TP, TA, LP, and LA. The retained factor explained 82.550% of the total variances. As shown in Table 15, the retained latent variable could capture all the features of the four different code similarity metrics. The detailed PCA analysis of code similarity are shown in B.
To validate the suitability of code similarity as a proxy for AI code adoption, we randomly selected 20 pairs of participants’ code—specifically, their original submissions before encountering the AI solution and their revised submissions after reviewing the AI-generated code. For each pair, we systematically evaluated two metrics: the similarity between their revised code and the AI solution, and the actual extent of AI code adoption.
Adoption levels were categorized as follows:
Full adoption: The participant entirely discarded their original code and adopted the AI-generated solution in its entirety, with no remnants of their initial implementation retained.
Partial adoption: The participant preserved portions of their original code while incorporating specific components of the AI-generated solution without full replication. The specific components adopted by participants are clarified in the Details column in Table LABEL:table:adoption-details. Examples include: additions of defensive programming elements (to enhance code robustness, such as incorporating if (!head) return head from the AI-generated solution to prevent null pointer errors), and adjustments to control structures (e.g., switching from recursive to iterative implementations to align with the AI solution).
No evidence of adoption: The revised code contains no recognizable elements derived from the AI solution.
As shown in Table LABEL:table:adoption-details, our analysis revealed a clear hierarchical pattern: revised code classified as full adoption showed significantly higher similarity to the AI solution than that with partial adoption, while partial adoption cases consistently exhibited greater similarity than those with no adoption. This ordered relationship confirms that code similarity effectively reflects the degree of AI code adoption, thereby validating its use as a reliable proxy measure.
{ThreePartTable}
Table 5: Adoption details of programming tasks
LeetCode Numbera Participant’s Prolific IDb Similarityc Adoption Level Details
83 6511 0.67 partial adoption defensive programming
83 60cd 0.81 full adoption
83 5f20 0.95 full adoption
83 5f0d -0.01 partial adoption defensive programming, control structure
98 657d 0.95 full adoption
98 6525 -0.11 partial adoption defensive programming
98 650b 0.95 full adoption
98 60ae -1.12 no evidence of adoption
1823 6611 0.95 full adoption
1823 65e1 -1.21 no evidence of adoption
1823 64eb 0.95 full adoption
1823 6129 -0.03 partial adoption defensive programming
32 654c 0.95 full adoption
32 627b 0.95 full adoption
32 5df6 -0.15 partial adoption defensive programming, control structure
32 5ba0 0.95 full adoption
383 6558 -1.74 no evidence of adoption
383 6122 0.95 full adoption
383 659e 0.95 full adoption
383 6609 -1.78 no evidence of adoption
Table 5: Adoption details of programming tasks (continued)
  • a
  • b
    The Participant’s Prolific ID column displays the first 4 digits of each participant’s Prolific ID. This ID serves as the unique identifier for participants on the Prolific platform.
  • c
    Negative values in the Similarity column exist because data was normalized.